# Spread and Spectral Complexity in Quantum Spin Chains: from Integrability to Chaos

### v1

@@ 1 Introduction
Remarkably, a one-to-one correspondence between a specific notion of quantum complexity and the length of the ERB was recently found in Rabinovici:2023yex within the context JT gravity and the double-scaled Sachdev–Ye–Kitaev (SYK) model SachdevYeModel ; KitaevTalks . The notion computed in the SYK model is Krylov state complexity, also known as spread complexity Balasubramanian:2022tpr . Krylov complexity (or simply K-complexity) was first was introduced in Parker:2018yvk as a measure of operator growth for operators undergoing Heisenberg time evolution in quantum many-body systems with local interactions. It was then extended to study the Schrödinger time evolution of states and measure their “spread” in a subspace of Hilbert space known as Kryolv space, a space spanned by successive applications of the quantum system’s Hamiltonian on an initial state. Spread and K-complexity have been studied in different contexts such as topological and quantum phase transitions Caputa:2022eye ; Afrasiar:2022efk ; Caputa:2022yju ; Pal:2023yik , operator growth Barbon:2019wsy ; Bhattacharjee:2022vlt , CFTs Dymarsky:2019elm ; Dymarsky:2021bjq ; Kundu:2023hbk , open quantum systems Bhattacharya:2022gbz ; Mohan:2023btr , and other related contexts Yates:2021asz ; Caputa:2021ori ; Patramanis:2021lkx ; Trigueros:2021rwj ; Rabinovici:2020ryf ; Rabinovici:2021qqt ; Rabinovici:2022beu ; Bhattacharjee:2022qjw ; Chattopadhyay:2023fob ; Bhattacharjee:2023dik ; Bhattacharjee:2022ave ; Takahashi:2023nkt ; Camargo:2022rnt ; Avdoshkin:2022xuw ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Iizuka:2023pov ; Caputa:2023vyr ; Fan:2023ohh ; Vasli:2023syq ; Gautam:2023bcm ; Iizuka:2023fba ; Huh:2023jxt ; Anegawa:2024wov ; Caputa:2024vrn ; Chen:2024imd ; Caputa:2024xkp . For a comprehensive and recent review of both spread and K-complexity, together with an extensive list of references, we refer the reader to Nandy:2024htc .
@@
We show that both spread and spectral complexity exhibit distinguishing features in chaotic spin systems which can be directly contrasted with their integrable counterparts. That is: the presence of a peak in spread complexity and a saturation of spectral complexity which is governed by the minimum value of the energy difference in the spectrum. These results can be seen as an effort in the same direction as studies performed for Krylov operator complexity Rabinovici:2020ryf ; Rabinovici:2021qqt ; Rabinovici:2022beu . In particular, for spread complexity, we verify that the characteristic peak is only manifest in the chaotic case, and that its late-time behaviour is model dependent. For the chaotic point of the Ising model, we find that the spread complexity saturates to a constant whose value is comparable with the one for the integrable case. In contrast, for the NNN deformation of the Heisenberg chain, we find that the saturation value of the integrable case lies below the value of the chaotic one. This is in direct analogy with the behaviour of Krylov (operator) complexity in the same models. Finally, for spectral complexity we confirm that its saturation value for the chaotic cases occurs much earlier than in their integrable counterparts. We also determine that this saturation occurs at a characteristic timescale in chaotic systems which is proportional to the inverse of the minimum energy difference in the Hamiltonian’s spectrum. This feature is explored for Hamiltonians belonging to different random matrix ensembles.
@@ 2.1 Spread Complexity
where $Z(\beta):=\textrm{tr}(e^{-\beta H})$ is the thermal partition function with inverse temperature $\beta$ and where $|n\rangle$ are energy eigenvalues $H|n\rangle=E_{n}|n\rangle$. We choose this initial state, since the peak structure of spread complexity was observed for it previously in the context of RMTs Balasubramanian:2022tpr ; Erdmenger:2023wjg .
@@ 3.1.2 The Heisenberg XXZ Model
For any values of $J$, $J_{zz}$ and $J_{c}$, this model displays parity symmetry, because $[H,\hat{\Pi}]=0$, with $\hat{\Pi}$ given by (21). Additionally, the model conserves the total spin in the $z$-direction, $M_{z}=\sum_{i=1}^{N}S_{i}^{z}$. For a detailed and pedagogical discussion of the symmetries of this model, we refer to Joel2013 .
@@ 3.2.1 Lanczos Coefficients and Peak
As mentioned previously, the NNN deformation of the Heisenberg Hamiltonian (23) also has two parity symmetry sectors, similarly to the mixed-field Ising model. The results for the spread complexity are shown in Fig. 7. Here, as shown in Fig. 5, the peak structure is also present in the chaotic case. At the same time, the late-time behaviour is also captured by (28) for the chaotic case.666 Nevertheless, it is not surprising that (28) fails to capture the late time behavior of the integrable case (blue data) depicted in Fig. 7. It has already been demonstrated Huh:2023jxt that the integrable models, such as LMG models, cannot adhere to (28). Note that the saturation value for the integrable case is noticeably smaller that the chaotic case, in contrast with the Ising chain (cfr. Fig. 5Balasubramanian:2022tpr ; Erdmenger:2023wjg .
@@ 3.2.2 A Comparison with the Spectral Form Factor
Our findings align with the argument in Erdmenger:2023wjg , suggesting that the transition time scales in both spread complexity and SFF are closely related. For example, we observe that the timescale of the peak in spread complexity and that of the plateau in the SFF factor both scale linearly with the system’s size $d$. See Fig. 16 for a detailed comparison.
@@
where the first equality is derived from Cotler:2016fpe , while the second is obtained through (28Erdmenger:2023wjg ; Rabinovici:2020ryf ; Rabinovici:2022beu . This is depicted as dashed lines in Fig. 9.888We make the plot of $1/d_{i}$ due to the potential invalidity of (28) for the integrable systems.
@@ 3.3.1 Early-Time Behavior and Characteristic Timescales
In this section we show the results for spectral complexity in spin chain models. We focus on two main aspects: the time-scale at which the saturation occurs and the early-time behaviour.
Regarding saturation, it is worth noting that for chaotic case, the saturation time-scale of spread complexity is of a comparable order to that of their integrable counterparts (refer to Sec. 2.1). However, there are examples in simple quantum systems, such as quantum billiards (Camargo:2023eev ), where the saturation time-scale of spectral complexity for chaotic systems is several orders of magnitude smaller than in their integrable counterparts. [^9] As advertised in the introduction, in Sec. 3.3.2 we provide an explanation for this disparity from the perspective of RMT at finite size $N$.
@@ 3.3.2 Lessons from Random Matrices: Minimum Energy Difference
Since the definition of spectral complexity (14) contains $\frac{1}{(E_{m}-E_{n})^{2}}$, the term with the smallest $|E_{m}-E_{n}|$ would contribute dominantly to the saturation value of $C_{S}(t)$. For the long-time average, we should take time longer than $\frac{\pi}{\Delta E_{\text{min}}}$, where $\Delta E_{\text{min}}:=\text{min}\{|E_{m}-E_{n}|\}$ is the minimum energy difference in quantum systems. Therefore, $\Delta E_{\text{min}}$ is an important inverse time scale for the saturation of spectral complexity. We study how $\Delta E_{\text{min}}$ depends on size $d$ of Hamiltonian matrix $H$.
@@ 4 Discussion and Conclusions
We also studied spectral complexity and provided further evidence to the observation made in Camargo:2023eev that its saturation value and timescale for the chaotic point of a given Hamiltonian that interpolates between integrable and chaotic, is orders of magnitude smaller than for the integrable point. By analyzing the properties of spectral complexity for random matrix theories (RMTs) of finite size $N$, we determined that the property of the system determining these features is the minimum of the energy difference in the spectrum. To the best of our knowledge, this is the first time that such a feature of the spectrum appeared in the context of quantum chaos.
Timescales of dynamical probes for quantum chaos.
Perhaps the most accepted definition of quantum chaos is based on the idea that chaotic systems behave like random matrices. The basic idea is that the energy levels of the system in each symmetry sector are strongly correlated, and their nearest-neighbor level spacing statistics follows a Wigner–Dyson distribution. By contrast, the energy levels of integrable systems are uncorrelated and their nearest-neighbor level spacing statistics typically follows a Poisson distribution. This definition is based on the Bohigas–Giannoni–Schmit (BGS) conjecture Bohigas:1983er , which states that certain fluctuation properties of the spectrum of time-reversal classically chaotic systems (K-systems) are well described by Gaussian orthogonal ensembles (GOE) of random matrices. This conjecture is still not proven but is supported by a vast number of numerical studies, with well-understood limitations in its validity. In the case of systems that do not have a classical limit, it is natural to use the random matrix behavior as a working definition of quantum chaos. This definition is very economical since it only depends on the Hamiltonian of the system. Nevertheless, in addition to studying the spectrum of the systems, which are static, exploring time-dependent quantities can also offer additional valuable insights for understanding manifestations of chaos at the quantum level. In particular, it is interesting to investigate how the chaotic properties of the Hamiltonian influence the dynamics of states and operators and what are the signatures of chaos for them. In Table. 1, including our findings on spread and spectral complexities, we summarize some well-studied probes of quantum chaos and their characteristic timescales.
@@
The standard time-dependent quantity in studies of quantum chaos is the Spectral Form Factor (SFF). In RMTs, SFF as a function of time displays a typical slope-dip-ramp-plateau behavior, characterized by some timescales. The same behavior is expected to occur in chaotic systems, while in integrable systems one expects a different behavior. Spread and spectral complexities provide a new window into this time dynamics.
@@
It is worth recall that these state-dependent (spread and spectral) complexities can probe random matrix behavior well when the complexities are associated with the TFD state. This happens because random matrix behavior is a statement about the full [^12] One should note, however, that random matrix behavior is normally found excluding the edges of the energy spectrum of the system, where the behavior is typically non-universal. Certain systems, however, display chaotic fluctuations through the entire spectrum Altland:2024ubs . (level-spacing) spectrum of the system. As such, in this regard, any quantity that aims at probing random matrix behavior would be beneficial to have information about the full spectrum of the system, and the TFD state accomplishes this. We believe that other states sufficiently similar to the TFD state are also capable of effectively probing random matrix behavior.

### v2

@@ 1 Introduction
Remarkably, a one-to-one correspondence between a specific notion of quantum complexity and the length of the ERB was recently found in Rabinovici:2023yex within the context JT gravity and the double-scaled Sachdev–Ye–Kitaev (SYK) model SachdevYeModel ; KitaevTalks . The notion computed in the SYK model is Krylov state complexity, also known as spread complexity Balasubramanian:2022tpr . Krylov complexity (or simply K-complexity) was first was introduced in Parker:2018yvk as a measure of operator growth for operators undergoing Heisenberg time evolution in quantum many-body systems with local interactions. It was then extended to study the Schrödinger time evolution of states and measure their “spread” in a subspace of Hilbert space known as Kryolv space, a space spanned by successive applications of the quantum system’s Hamiltonian on an initial state. Spread and K-complexity have been studied in different contexts such as topological and quantum phase transitions Caputa:2022eye ; Afrasiar:2022efk ; Caputa:2022yju ; Pal:2023yik , operator growth Barbon:2019wsy ; Bhattacharjee:2022vlt , CFTs Dymarsky:2019elm ; Dymarsky:2021bjq ; Kundu:2023hbk ; Malvimat:2024vhr , open quantum systems Bhattacharya:2022gbz ; Bhattacharjee:2022lzy ; Mohan:2023btr ; Bhattacharya:2023zqt ; Bhattacharjee:2023uwx ; Carolan:2024wov , and other related contexts Yates:2021asz ; Caputa:2021ori ; Patramanis:2021lkx ; Trigueros:2021rwj ; Rabinovici:2020ryf ; Rabinovici:2021qqt ; Rabinovici:2022beu ; Bhattacharya:2023xjx ; Bhattacharjee:2022qjw ; Chattopadhyay:2023fob ; Bhattacharjee:2023dik ; Bhattacharjee:2022ave ; Takahashi:2023nkt ; Camargo:2022rnt ; Avdoshkin:2022xuw ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Iizuka:2023pov ; Caputa:2023vyr ; Fan:2023ohh ; Vasli:2023syq ; Gautam:2023bcm ; Iizuka:2023fba ; Huh:2023jxt ; Anegawa:2024wov ; Caputa:2024vrn ; Chen:2024imd ; Caputa:2024xkp ; Chattopadhyay:2024pdj ; Nandy:2023brt ; Aguilar-Gutierrez:2023nyk ; Aguilar-Gutierrez:2024nau . For a comprehensive and recent review of both spread and K-complexity, together with an extensive list of references, we refer the reader to Nandy:2024htc .
@@
We show that both spread and spectral complexity exhibit distinguishing features in chaotic spin systems which can be directly contrasted with their integrable counterparts. That is: the presence of a peak in spread complexity and a saturation of spectral complexity which is governed by the minimum value of the energy difference in the spectrum. These results can be seen as an effort in the same direction as studies performed for Krylov operator complexity Rabinovici:2020ryf ; Rabinovici:2021qqt ; Rabinovici:2022beu ; Bhattacharya:2023xjx . In particular, for spread complexity, we verify that the characteristic peak is only manifest in the chaotic case, and that its late-time behaviour is model dependent. For the chaotic point of the Ising model, we find that the spread complexity saturates to a constant whose value is comparable with the one for the integrable case. In contrast, for the NNN deformation of the Heisenberg chain, we find that the saturation value of the integrable case lies below the value of the chaotic one. This is in direct analogy with the behaviour of Krylov (operator) complexity in the same models. Finally, for spectral complexity we confirm that its saturation value for the chaotic cases occurs much earlier than in their integrable counterparts. We also determine that this saturation occurs at a characteristic timescale in chaotic systems which is proportional to the inverse of the minimum energy difference in the Hamiltonian’s spectrum. This feature is explored for Hamiltonians belonging to different random matrix ensembles.
@@ 2.1 Spread Complexity
where $Z(\beta):=\textrm{tr}(e^{-\beta H})$ is the thermal partition function with inverse temperature $\beta$ and where $|n\rangle$ are energy eigenvalues $H|n\rangle=E_{n}|n\rangle$. [^4] In the presence of symmetries, we construct the TFD state separately for each symmetry sector of the system. We choose this initial state, since the peak structure of spread complexity was observed for it previously in the context of RMTs Balasubramanian:2022tpr ; Erdmenger:2023wjg .
@@ 3.1.2 The Heisenberg XXZ Model
For any values of $J$, $J_{zz}$ and $J_{c}$, this model displays parity symmetry, because $[H,\hat{\Pi}]=0$, with $\hat{\Pi}$ given by (21). Additionally, the model conserves the total spin in the $z$-direction, $M_{z}=\sum_{i=1}^{N}S_{i}^{z}$. For a detailed and pedagogical discussion of the symmetries of this model, we refer the reader to Joel2013 .
@@ 3.2.1 Lanczos Coefficients and Peak
As mentioned previously, the NNN deformation of the Heisenberg Hamiltonian (23) also has two parity symmetry sectors, similarly to the mixed-field Ising model. The results for the spread complexity are shown in Fig. 7. Here, as shown in Fig. 5, the peak structure is also present in the chaotic case. At the same time, the late-time behaviour is also captured by (28) for the chaotic case.777 Nevertheless, it is not surprising that (28) fails to capture the late time behavior of the integrable case (blue data) depicted in Fig. 7. It has already been demonstrated Huh:2023jxt that the integrable models, such as LMG models, cannot adhere to (28). In the derivation of (28), it is assumed that there is no energy degeneracy. In numerical calculations on integrable systems, approximate energy degeneracy within numerical precision may occur due to the Poisson distribution of the level spacing. For this reason, within the numerical precision and time scale of our numerical calculations, (28) in the integrable case does not seem to hold in Fig. 7. Note that the saturation value for the integrable case is noticeably smaller that the chaotic case, in contrast with the Ising chain (cfr. Fig. 5Balasubramanian:2022tpr ; Erdmenger:2023wjg .
@@ 3.2.2 A Comparison with the Spectral Form Factor
Our findings align with the argument in Erdmenger:2023wjg , suggesting that the transition time scales in both spread complexity and SFF are closely related. For example, we observe that the timescale of the peak in spread complexity and that of the plateau in the SFF both scale linearly with the system’s size $d$. See Fig. 16 for a detailed comparison.
@@
where the first equality is derived from Cotler:2016fpe , while the second is obtained through (28Erdmenger:2023wjg ; Rabinovici:2020ryf ; Rabinovici:2022beu . This is depicted as dashed lines in Fig. 9.999We make the plot of $1/d_{i}$ since  (28) does not hold for systems exhibiting degeneracy within the energy spectrum, as discussed in footnote 7.
@@ 3.3.1 Early-Time Behavior and Characteristic Timescales
In this section we show the results for spectral complexity in spin chain models. We focus on two main aspects: the time-scale at which the saturation occurs and the early-time behavior.
Regarding saturation, it is worth noting that for the chaotic case, the saturation time-scale of spread complexity is of a comparable order to that of their integrable counterparts (refer to Sec. 2.1). However, there are examples in simple quantum systems, such as quantum billiards (Camargo:2023eev ), where the saturation time-scale of spectral complexity for chaotic systems is several orders of magnitude smaller than in their integrable counterparts. [^10] As advertised in the introduction, in Sec. 3.3.2 we provide an explanation for this disparity from the perspective of RMT at finite size $N$.
@@ 3.3.2 Lessons from Random Matrices: Minimum Energy Difference
Since the definition of spectral complexity (14) contains $\frac{1}{(E_{m}-E_{n})^{2}}$, the term with the smallest $|E_{m}-E_{n}|$ would contribute dominantly to the saturation value of $C_{S}(t)$. For the long-time average, we should take a longer time than $\frac{\pi}{\Delta E_{\text{min}}}$, where $\Delta E_{\text{min}}:=\text{min}\{|E_{m}-E_{n}|\}$ is the minimum energy difference in quantum systems. Therefore, $\Delta E_{\text{min}}$ is an important inverse time scale for the saturation of spectral complexity. We study how $\Delta E_{\text{min}}$ depends on size $d$ of Hamiltonian matrix $H$.
@@ 4 Discussion and Conclusions
It is worth pointing out that the time-dependent properties of other quantum complexity measures, such as time-evolution of Nielsen complexity, have also been studied in the context of integrable and chaotic systems. In particular, a bound on Nielsen complexity has been efficiently estimated for bi-invariant metrics using lattice cryptography techniques Craps:2022ese ; Craps:2023rur . Building on these results, authors in Craps:2023ivc furthermore uncovered a relation between the time average of spread complexity and a bound on Nielsen complexity. It would be interesting to continue to explore the relation between these two complexity measures and understand whether there exist bounds to the growth rate of spread complexity akin to Hornedal:2022pkc , that could help understand the physics behind its saturation value, such as its state dependence.
We also studied spectral complexity and provided further evidence to the observation made in Camargo:2023eev that its saturation value and timescale for the chaotic point of a given Hamiltonian that interpolates between integrable and chaotic, is orders of magnitude smaller than for the integrable point. By analyzing the properties of spectral complexity for random matrix theories (RMTs) of finite size $d=2^{N}$, we determined that the property of the system determining these features is the minimum of the energy difference in the spectrum. To the best of our knowledge, this is the first time that such a feature of the spectrum appeared in the context of quantum chaos.
Timescales of dynamical probes of quantum chaos.
Perhaps the most accepted definition of quantum chaos is based on the idea that Hamiltonians of chaotic systems behave like random matrices. The basic idea is that the energy levels of the system in each symmetry sector are strongly correlated, and their nearest-neighbor level spacing statistics follows a Wigner–Dyson distribution. By contrast, the energy levels of integrable systems are uncorrelated and their nearest-neighbor level spacing statistics typically follows a Poisson distribution. This definition is based on the Bohigas–Giannoni–Schmit (BGS) conjecture Bohigas:1983er , which states that certain fluctuation properties of the spectrum of time-reversal classically chaotic systems (K-systems) are well described by Gaussian orthogonal ensembles (GOE) of random matrices. This conjecture is still not proven but is supported by a vast number of numerical studies, with well-understood limitations in its validity. In the case of systems that do not have a classical limit, it is natural to use the random matrix behavior as a working definition of quantum chaos. This definition is very economical since it only depends on the Hamiltonian of the system. Nevertheless, in addition to studying the spectrum of the systems, which are static, exploring time-dependent quantities can also offer additional valuable insights for understanding manifestations of chaos at the quantum level. In particular, it is interesting to investigate how the chaotic properties of the Hamiltonian influence the dynamics of states and operators and what are the signatures of chaos for them. In Table. 1, including our findings on spread and spectral complexities, we summarize some well-studied probes of quantum chaos and their characteristic timescales.
@@
The standard time-dependent quantity in studies of quantum chaos is the Spectral Form Factor (SFF). In RMTs, SFF as a function of time displays a typical slope-dip-ramp-plateau behavior, characterized by certain timescales. The same behavior is expected to occur in chaotic systems, while in integrable systems one expects a different behavior. Spread and spectral complexities provide a new window into this time dynamics.
@@
It is worth recalling that these state-dependent (spread and spectral) complexities can probe random matrix behavior well when the complexities are associated with the TFD state. This happens because random matrix behavior is a statement about the full [^13] One should note, however, that random matrix behavior is normally found excluding the edges of the energy spectrum of the system, where the behavior is typically non-universal. Certain systems, however, display chaotic fluctuations through the entire spectrum Altland:2024ubs . (level-spacing) spectrum of the system. As such, in this regard, any quantity that aims at probing random matrix behavior would benefit from having information about the full spectrum of the system, and the TFD state accomplishes this. We believe that other states sufficiently similar to the TFD state are also capable of effectively probing random matrix behavior.